# Transient selection of competing order in frustrated spin Peierls systems after a quench

### v1

@@ Abstract
We now compare the evolution of non-collinear magnetic order for different phonon relaxation time ratios $\tau$ (Fig. 4). The overall magnitude of $S^{y}(q=0)$ and $\xi^{y}$ increases with $\tau$. The correlation length $\xi^{y}$ for $\tau=150$ and $\tau=500$ both show an initial approximate $t^{1/2}$ growth, followed by a slower growth after the spin reorientation time $t_{c}$. The data for $\tau=2000$ only shows the approximate $t^{1/2}$ growth within the simulated time window.
To conclude, we have demonstrated that a frustrated spin-Peierls system can exhibit a transient, competing magnetic order owing to the nonequilibrium, incoherent phonon fluctuations. While the analysis is carried out for a minimal model derived from the BCC antiferromagnet, the idea applies to systems with other lattice geometries given its simplicity. Potential experimental platforms include magnets CuFeO2 [36, 37, 38, 39, 40], CuMnO2 [41, 42] and NaMnO2 [43, 44, 45], which feature geometric frustration and strong magnetoelastic coupling, as well as engineered systems such as buckled colloids [46].
Viewed from a broader perspective, our work points to a couple of directions for further research. Our setup is mapped to a peculiar phase ordering process, namely an XY model whose easy-axis anisotropy changes sign in time. An analysis of this case will shed light on how the spatial correlation of the transient order evolves in time. More importantly, the impact of nonequilibrium fluctuations on competing phases unveiled in this work may be fruitfully explored in the context of intertwined and vestigial orders [47, 48].
@@ Appendix A Construction of the Landau free energy
Half translation along $[111]$ $\theta_{A}\leftrightarrow\theta_{B}$
@@
We recall that, microscopically, the spin-phonon coupling is of the form $Q\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, After coarse-graining, the spin bilinear $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ would give rise to terms of the form $\cos(m\theta_{A}+n\theta_{B})$ or $\sin(m\theta_{A}+n\theta_{B})$. The spin rotational symmetry rules out terms with $m\neq-n$. We thus only need to consider $\cos(m\theta_{A}-m\theta_{B})$ and $\sin(m\theta_{A}-m\theta_{B})$. The half translation symmetry rules out the sine harmonics. As for the cosine harmonics, they changes sign under the inversion when $m$ is odd, and remains invariant when $m$ is even. Since we seek terms that would break the lattice symmetry, the relevant, lowest order harmonics is $\cos(\theta_{A}-\theta_{B})$. We thus obtain the spin-phonon coupling term:
@@
It is easy to verify that $\cos(\theta_{A}-\theta_{B})$ carries the $A_{2u}$ irreducible representation of the $O_{h}$ point group. Therefore, the phonon modes $Q$ must also carry the same representation.
@@ Appendix B Additional data for the structure factor
Fig. 5a&c illustrate the time evolution of the structure factors $S^{x,y}(\bm{q})$ along the $(100)$ direction. We estimate the location of the half width at half maximum, $q_{0}$, by a simple linear interpolation (triangular markers). Because the Bragg peaks become quite narrow at late time, only two momentum points are useable for the presented system size $L=120$ (data for $t=600$ and $t=900$). This results in a larger degrees of uncertainty in determining $q_{0}$, which, in turn, translates to larger uncertainty in estimating the correlation length $\xi^{x,y}$.

### v2

@@ Abstract
We now compare the evolution of non-collinear magnetic order for different phonon relaxation time ratios $\tau$ (Fig. 4). The overall magnitude of $S^{y}(q=0)$ and $\xi^{y}$ increases with $\tau$. The correlation length $\xi^{y}$ for $\tau=150$ and $\tau=500$ both show an initial approximate $\widetilde{t}^{1/2}$ growth, followed by a slower growth after the spin reorientation time $t_{c}$. The data for $\tau=2000$ only show the approximate $\widetilde{t}^{1/2}$ growth within the simulated time window.
To conclude, we have demonstrated that a frustrated spin-Peierls system can exhibit a transient, competing magnetic order owing to the nonequilibrium, incoherent phonon fluctuations. While the analysis is carried out for a BCC lattice model for simplicity, whose material incarnations are rare, we expect the mechanism applies to systems with other lattice geometries. Aside from its square lattice cousin [46, 47], preliminary analysis of triangular and pyrochlore lattice models, motivated by delafossites [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] and chromate spinels [59, 40, 60, 61, 62, 63, 64, 37], show that the non-equilibrium phonon fluctuations favor magnetic orders that are distinct from those by lattice distortion [35].
The order by quantum/thermal disorder can be incorporated into our model by adding a term $\Delta\cos^{2}(\theta_{A}-\theta_{B})$ [35], which suppresses the non-collinear state. In the quantum case, $\Delta\propto 1/S$, where $S$ is the quantum spin number. It is likely modest in delafossites and chromate spinels for their large $S$. In the thermal case, $\Delta\propto\widetilde{T}_{f}$, tunable with experimental conditions. The spatial correlation of phonon fluctuations, omitted in the present analysis, do not alter the transient selection of competing order provided that the correlation is short-ranged.
Our work points to a couple of directions for further research. The setup is mapped to a peculiar phase ordering process, namely an XY model whose easy-axis anisotropy changes sign in time. An analysis of this case will shed light on how the spatial correlation of the transient order evolves in time. Our model can also be extended to the case of underdamped phonons. Preliminary numerical simulation shows, after a quench, they produce qualitative similar behavior as underdamped phonons in the slow relaxation regime [35]. The impact of nonequilibrium fluctuations on competing phases unveiled in this work may be fruitfully explored in the context of intertwined and vestigial orders [65, 66, 24, 67].
Acknowledgements.
This work is supported by the National Key R&D Program of China (Grant No.2022YFA1403800, 2024YFA1408700), National Natural Science Foundation of China (Grant No.12250008, 11974396, and 12188101), and by the CAS Project for Young Scientists in Basic Research (Grant No. YSBR-059).
The data that support the findings of this article are openly available [68].
@@ Appendix A Construction of the Landau free energy
Half translation along $[111]$ $\theta_{B}\to\theta_{A}$, $\theta_{A}\to\theta_{B}+\pi$
@@
We recall that, microscopically, the spin-phonon coupling is of the form $Q\mathbf{S}_{i}\cdot\mathbf{S}_{j}$. In this work, we consider the so-called bond phonon model for the sake of simplicity: we assume that the displacement in the anions that mediate the super-exchange modulates the strength of exchange interaction, and that the modulation on these bonds are independent from each other. The modulation pattern shown in the main text is uniform within each simple cubic sub-lattice but odd under the half-translation along $[111]$. This pattern corresponds to the condensation of a phonon mode $Q$ that is at the high symmetry point $H=[100]$ of the first Brillouin zone. The symmetry of this phonon mode is classified by the little co-group $O_{h}$. It is easy to see that this mode carries irreducible representation $A_{2u}$.
After coarse-graining, the spin bilinear $\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ would give rise to terms of the form $\cos(m\theta_{A}+n\theta_{B})$ or $\sin(m\theta_{A}+n\theta_{B})$. The spin rotational symmetry rules out terms with $m\neq-n$. We thus only need to consider $\cos(m\theta_{A}-m\theta_{B})$ and $\sin(m\theta_{A}-m\theta_{B})$. The symmetry requires that this term is odd under half-translation and carries the $A_{2u}$ representation. The half translation symmetry rules out the sine harmonics. As for the cosine harmonics, the lowest order harmonics is $\cos(\theta_{A}-\theta_{B})$, and shows the same transformation properties as $Q$. We thus obtain the spin-phonon coupling term:
@@
We discuss briefly on the connection of the effective model parameters to microscopic model parameters. To this end, we consider a microscopic Hamiltonian:
$H=J_{1}\sum_{\langle ij\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+J_{2}\sum_{\langle\langle ij\rangle\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}-J^{\prime}\sum_{i\in A}\sum_{n}Q_{i}\eta_{n}\mathbf{S}_{i}\cdot\mathbf{S}_{i+n}+\frac{k}{2}\sum_{i\in A}Q^{2}_{i}.$ (7)
Here, the first and second terms sum over the nearest and second neighbor pairs in the BCC lattice. In the third term, we sum over all bonds $n$ that emanate from an $A$ site, namely $n=(\pm a/2,\pm a/2,\pm a/2)$, where $a$ is the lattice constant. $\eta_{n}$ is a form factor:
$\eta_{n}=\left\{\begin{array}[]{cc}1&n=(a/2,a/2,a/2),(a/2,-a/2,-a/2),(-a/2,a/2,-a/2),(-a/2,-a/2,a/2)\\ -1&n=(-a/2,-a/2,-a/2),(-a/2,a/2,a/2),(a/2,-a/2,a/2),(a/2,a/2,-a/2)\end{array}\right.$. (10)
Note this model does not break the symmetry between the A and B sub-lattices: the third sum can be rewritten in terms of B sublattices with $Q\to-Q$.
We proceed to coarse-grain the above model by using the gradient expansion. We postulate:
$\mathbf{S}_{i\in A}=\mathbf{N}_{A}(r_{i})e^{iK\cdot r_{i}};\quad\mathbf{S}_{i\in B}=\mathbf{N}_{B}(r_{i})e^{iK\cdot(r_{i}-n_{0})};\quad Q_{i}=Q(r_{i}),$ (11)
where the right hand sides are smooth functions of the coordinates. $n_{0}=(a/2,a/2,a/2)$. $K=(\pi,\pi,\pi)/a$ is the characteristic wave vector of the Néel order. To the quadratic order in lattice constant $a$, we have:
$H=\int\frac{d^{3}x}{a^{3}}\{\frac{J_{2}a^{2}}{2}[(\nabla\mathbf{N}_{A})^{2}+(\nabla\mathbf{N}_{B})^{2}]-8J^{\prime}Q\mathbf{N}_{A}\cdot\mathbf{N}_{B}+\frac{k}{2}Q^{2}\}.$ (12)
Setting $\mathbf{N}_{A}=(\cos\theta_{A},\sin\theta_{A})$ and $\mathbf{N}_{B}=(\cos\theta_{B},\sin\theta_{B})$, we obtain the expression of Landau free energy in the main text:
$H=\int d^{3}x\{\frac{K}{2}[(\nabla\theta_{A})^{2}+(\nabla\theta_{B})^{2}]-\lambda Q\cos(\theta_{A}-\theta_{B})+\frac{\epsilon}{2}Q^{2}\},$ (13)
with
$K=\frac{J_{2}}{a};\quad\lambda=\frac{8J^{\prime}}{a^{3}};\quad\epsilon=\frac{k}{a^{3}}.$ (14)
Interestingly, we see that $\kappa=0$ to the leading order of gradient expansion. It can be generated by integrating out short-ranged fluctuations, or by coarse-graining multi-spin interactions. We also see that $J_{1}$ term does not contribute at this order (it contributes to the third order in the gradient expansion).
@@ Appendix B Additional data for the structure factor
Fig. 5a&c illustrate the time evolution of the structure factors $S^{x,y}(\bm{q})$ along the $(100)$ direction. We estimate the location of the half width at half maximum, $q_{0}$, by a simple linear interpolation (triangular markers). Because the Bragg peaks become quite narrow at late time, only two momentum points are useable for the presented system size $L=120$ (data for $\widetilde{t}=600$ and $\widetilde{t}=900$). This results in a larger degrees of uncertainty in determining $q_{0}$, which, in turn, translates to larger uncertainty in estimating the correlation length $\xi^{x,y}$.